Find the linear approximation to f(x,y,z)=xy/z at the point (2,2,−1):

f(x,y,z)≈
1
Expert's answer
2020-10-14T16:58:58-0400

f(x,y,z)=xy/z,

f(2,2,-1)=2*2/(-1)=4,

f`x= y/z , f`x(2,2,-1)=2/(-1) =-2,

f`y= x/z , f`y(2,2,-1)=2/(-1) =-2,

f`z= -xy/z2 , f`z(2,2,-1)=-2*2/(-1)2=-4,


f(x,y,z) = f(2,2,-1) + f`x(2,2,-1)*(x-2)+ f`y(2,2,-1)*(y-2)+ f`z(2,2,-1)*(z+1),

f(x,y,z) = 4+(-2)(x-2)+(-2)(y-2)+(-4)(z+1) = 8-2x-2y-4z,


f(x,y,z)=4-x-y-2z .